Find the general solution frac{dy}{dx}= frac{2y^2+3xy}{x^2}
1
Expert's answer
2019-10-11T10:07:32-0400

This is a Bernoulli differential equation.

"\frac{dy}{dx}= \frac{2y^2}{x^2}+\frac{3y}{x}"

Let's divide both sides by "\,-y^2" "\,"

"-\frac{dy}{y^2dx}= -\frac{2}{x^2}-\frac{3}{xy}"

Using the substitution "v=y^{-1}" , "v'=-\frac{y'}{y^2}" will lead us to "\,"

"v'=-\frac{2}{x^2}-\frac{3v}{x}"

This is a first-order linear equation which can be solved using an integrating factor "\,"

"M=e^{\int\frac{3}{x}dx}=e^{3\ln x}=x^3"

Multiplying both sides by M we get "\,"

"x^3v'+3x^2v=-2x\\"

"(x^3v)'=-2x\\"

"x^3v=-x^2+C"

"v=\frac{C-x^2}{x^3}\\"

Hence "\,"

"y=\frac{1}{v}=\frac{x^3}{C-x^2}"